# FuSeFL: Fully Secure and Scalable Cross-Silo Federated Learning

### v2

@@ VII-F Robustness Under Threat Model
Protection Against Unauthorized Observation. Adversaries may attempt to extract sensitive information by intercepting communication between clients or by observing model parameters during training. In FuSeFL, all private values, including client data, model parameters, and gradients, are secret-shared using additive masking. These shares appear statistically indistinguishable from random noise, preventing inference even in the case of eavesdropping.
To securely perform training, FuSeFL employs the AriaNN protocol [25], which guarantees confidentiality through function sharing and additive masking. During secure training, no single party observes plaintext data or model parameters. Additionally, each client retains one share of its data locally and never reveals its complete input to its training partner, preserving local privacy. These guarantees are backed by established cryptographic protocols validated in prior literature [22, 43, 45, 39, 40].
Protection Against Model Theft. Model confidentiality is a critical requirement in federated settings, particularly when the model represents proprietary intellectual property. In FuSeFL, the global model is secret-shared between two non-colluding servers and then transmitted to clients as encrypted shares. Each client receives only one share and cannot reconstruct the model independently. Even if an adversary intercepts a share during transmission, it reveals no meaningful information.
Furthermore, aggregation servers do not have access to the full model updates; they operate only on shares during the FedAvg procedure. These protections collectively ensure that no party, client or server, can exfiltrate the full model under the semi-honest assumption. FuSeFL thereby mitigates the risk of model theft during both training and communication phases.
@@ VIII Related Work
To protect privacy in FL, several cryptographic and statistical methods have been proposed. Homomorphic encryption (HE) allows computation over encrypted data but is computationally intensive, especially in cross-silo settings with large-scale datasets [3, 4]. Differential privacy (DP) introduces calibrated noise to gradients or updates to provide formal privacy guarantees [5, 6], but this often comes at the cost of reduced model accuracy. Secure multi-party computation (MPC), in contrast, offers strong privacy guarantees without degrading accuracy, though it typically incurs high communication costs.
MPC-based approaches have gained traction in secure FL. SecureML [22], SecureNN [23], Falcon [24], and AriaNN [25] propose secure training protocols using various cryptographic primitives, including additive secret sharing [40, 39], function secret sharing [43, 45], and Yao’s circuits [49]. These works typically offload both data and models to trusted third-party servers for secure computation. However, they assume trust in infrastructure and require clients to fully transmit their data shares, increasing bandwidth demands and collusion risks.
WW-FL [20] extends this idea by introducing a cluster-based MPC training framework. Clients are grouped into clusters and offload secret-shared data to server pools, which perform collaborative training. While this improves server-side load distribution, it still assumes trusted execution on third-party servers and leads to high communication overhead as the number of clients or cluster sizes increase. Additionally, WW-FL relies on a client-side trusted zone for safely outsourcing data, which may not be feasible in practical deployments.
@@
Finally, hierarchical FL [50, 51, 20] reduces communication costs by introducing intermediate aggregators for groups of clients. This approach improves scalability and bandwidth usage but does not inherently address the confidentiality of models or data during training.
@@ IX Conclusion
In this work, we introduced FuSeFL, a fully secure and scalable federated learning framework tailored for cross-silo environments. By decentralizing training across client pairs using lightweight MPC and assigning a minimal aggregation role to the servers, FuSeFL eliminates central bottlenecks and ensures complete confidentiality of client data, model parameters, and training updates. Our design not only mitigates key inference threats, including model inversion and gradient leakage, but also significantly reduces communication latency and server resource demands. Empirical evaluations show that FuSeFL achieves strong accuracy and performance improvements over existing secure FL solutions, establishing it as a practical and robust alternative for secure collaborative learning in regulated domains.

### v3

@@ VII-E Robustness Under Threat Model
Protection Against Unauthorized Observation. Adversaries may attempt to extract sensitive information by intercepting communication between clients or by observing model parameters during training. In FuSeFL, all private values, including client data, model parameters, and gradients, are secret-shared using additive masking. These shares appear statistically indistinguishable from random noise, preventing inference even in the case of eavesdropping. To securely perform training, FuSeFL employs the AriaNN protocol [22], which guarantees confidentiality through function sharing and additive masking. During secure training, no single party observes plaintext data or model parameters. Additionally, each client retains one share of its data locally and never reveals its complete input to its training partner, preserving local privacy. These guarantees are backed by established cryptographic protocols validated in prior literature [19, 51, 53, 47, 48].
Protection Against Model Theft. Model confidentiality is a critical requirement in federated settings, particularly when the model represents proprietary intellectual property. In FuSeFL, the global model is secret-shared between two non-colluding servers (at least one of which is trusted) and then transmitted to clients as encrypted shares. Each client receives only one share and cannot reconstruct the model independently. Even if an adversary intercepts a share during transmission, it reveals no meaningful information. In addition, the design of FuSeFL is backed up by the idea of mix networks, where the trusted server acts as a mix node to ensure that clients are grouped anonymously and cannot collude. Furthermore, aggregation servers do not have access to the full model updates; they operate only on shares during the FedAvg procedure. These protections collectively ensure that no party, client or server, can exfiltrate the full model under the semi-honest assumption. FuSeFL thereby mitigates the risk of model theft during both training and communication phases.
@@ VIII Related Work
To protect privacy in FL, several cryptographic and statistical methods have been proposed. HE allows computation over encrypted data but is computationally intensive [3, 4]. DP introduces calibrated noise to gradients or updates to provide formal privacy guarantees [5, 6], but this often comes at the cost of reduced model accuracy. MPC, in contrast, offers strong privacy guarantees without degrading accuracy, though it typically incurs high communication costs. MPC-based approaches have gained traction in secure FL. SecureML [19], SecureNN [20], Falcon [21], and AriaNN [22] propose secure training protocols using various cryptographic primitives, including additive secret sharing [48, 47], function secret sharing [51, 53], and Yao’s circuits [58]. These works typically offload both data and models to trusted third-party servers for secure computation. However, they assume trust in infrastructure and require clients to fully transmit their data shares, increasing bandwidth demands and collusion risks. WW-FL [18] extends this idea by introducing a cluster-based MPC training framework. Clients are grouped into clusters and offload secret-shared data to server pools, which perform collaborative training. While this improves server-side load distribution, it still assumes trusted execution on third-party servers and leads to high communication overhead as the number of clients or cluster sizes increase. Additionally, WW-FL relies on a client-side trusted zone for safely outsourcing data, which may not be feasible in practical deployments.
@@ IX Conclusion
In this work, we introduced FuSeFL, a fully secure and scalable federated learning framework tailored for cross-silo environments. By decentralizing training across client pairs using lightweight MPC and assigning a minimal aggregation role to the servers, FuSeFL eliminates central bottlenecks and ensures complete confidentiality of client data, model parameters, and training updates. Our design not only mitigates key inference threats, including model inversion and gradient leakage, but also significantly reduces training time and server resource demands. Our evaluations show that FuSeFL achieves performance improvements over existing secure FL solutions, establishing it as a practical and robust alternative for secure collaborative learning in regulated domains.